# Faithful Group Shapley Value

### v1

@@ Results.
Figure 2 summarizes the average performance over $30$ replications for $n\in\{64,128,256\}$. The top row shows that our method consistently achieves the lowest AUCC across all problem sizes $n$. The bottom row suggests that, while using the same utility evaluation budget, our method is among the fastest. In contrast, baselines such as Group Testing, KernelSHAP, and Unbiased KernelSHAP incur substantial higher computational overhead, due to internal optimization processes. Overall, our method achieves both faster convergence and improved computational efficiency, which demonstrates the benefits of directly estimating the group value rather than aggregating individual Shapley estimates.
@@ 4.2 Application to faithful copyright attribution in generative AI
Group data valuation is important for fairly compensating copyright holders whose data are used to train generative AI models. Existing approaches, such as the Shapley Royalty Share (SRS) proposed by [35], adopt the GaI approach based on GSV. Given a partition of the training data into $K$ disjoint groups $S_{1},\dots,S_{K}$, SRS is defined as $\operatorname{SRS}(S_{k};x_{\mathrm{gen}}):=\frac{\operatorname{GSV}(S_{k};x_{\mathrm{gen}})}{\sum_{j=1}^{K}\operatorname{GSV}(S_{j};x_{\mathrm{gen}})}$. However, as discussed in Section 2, GSV is prone to the shell company attack. To address this vulnerability, we propose the Faithful Shapley Royalty Share (FSRS) that replaces GSV with our $\operatorname{FGSV}$ to faithfully reward individuals in a group by their contributions, not by tactical grouping strategies:
@@
Following [35], we fine-tune Stable Diffusion v1.4 [28] using Low-Rank Adaptation (LoRA; [11]) on four brand logos from FlickrLogo-27 [15]. The utility $U(\cdot;x^{(\mathrm{gen})})$ is the average log-likelihood of generating $20$ brand-specific images $x^{(\mathrm{gen})}$ using the prompt “A logo by [brand name]” (see example images in Panel (a) of Figure 3). We compare SRS and FSRS under two grouping scenarios: (1) $30$ images from each brand form a single group, and (2) the Google and Sprite datasets are each split into two subgroups (20/10 images), launching a shell company attack. Importantly, the total data per brand remains unchanged across two scenarios.
Figure 3: Comparison of SRS and FSRS for copyright attribution. (a) Example images generated using brand prompts. (b) Shapley Royalty Share (SRS, [35]) based on GSV. (c) Faithful SRS (FSRS, our method) based on $\operatorname{FGSV}$. Blue bars: valuation under Scenario 1 (single group per brand); orange bars: valuation under Scenario 2 (Google/Sprite data each split into size-20/10 subgroups, colored in dark and light orange).
@@ 4.3 Application to faithful explainable AI
Group data valuation is also a crucial tool in explainable AI, providing interpretable summaries of data contributions at the group level— particularly in contexts where groups correspond to socially or scientifically significant categories. For example, [6] used GSV to quantify the contributions of data from different demographic groups to patient readmission prediction accuracy [31]. However, GSV’s sensitivity to the choice of grouping can cause group values to fluctuate dramatically under different data partitions, leading to inconsistent interpretations. To address this problem, we recommend practitioners to use FGSV and empirically demonstrate that it produces more consistent and reliable interpretations across a variety of grouping configurations.
@@
We conduct our experiment on the Diabetes dataset [5], which contains 442 individuals, each described by 10 demographic and health-related features (e.g., sex, age, and BMI). The task is to predict the progression of diabetes one year after baseline. We construct 7 grouping schemes, partitioned by all non-empty subsets of the variables sex, age, and BMI, excluding the trivial no-partition case. For the continuous variables age and BMI, we discretize each into three quantile-based categories. Overall, we will have between 2 to 18 total groups. For each grouping scheme, we compute GSV exactly and estimate FGSV via 30 Monte Carlo replications. Our predictive model is ridge regression, and we measure utility as the negative mean squared error on a held-out test set, with the null utility set to the variance of the test responses. Thus, higher group valuations indicate greater contributions to the model’s predictive accuracy. To compare category-level contributions across different grouping schemes, we aggregate group values as follows: for a given variable (e.g., sex), we sum the (F)GSVs of all groups that include that category. This allows for consistent cross-scheme comparisons, where one would expect a category to receive stable valuations after aggregation across different grouping strategies if the valuation method is robust.
Figure 4: Comparison of GSV (top row) and FGSV (bottom row) in a regression task for explainable AI. Each column aggregates category-level values for a specific variable: sex (left), age (middle), and BMI (right). Shaded areas represent ±$\pm 1$ standard deviation across 30 replications.
@@ 5 Conclusion and discussion
Beyond the shell company attack, there exists another unfair competition strategy, namely, the copier attack, in which a group may steal valuation from other groups by duplicating their high-value data points. While a principled defense against this attack remains an open problem, a straightforward ad-hoc remedy is to pre-process the dataset to detect and remove such duplicated entries before applying our FGSV method.

### v2

@@ Results.
Figure 2 summarizes the average performance over $30$ replications for $n\in\{64,128,256\}$. The top two rows together indicate that our method shows overall superior performance across all problem sizes $n$, exhibiting both the lowest average AUCC and ARE. The bottom row suggests that, while using the same utility evaluation budget, our method is among the fastest. In contrast, baselines such as Group Testing, KernelSHAP, and Unbiased KernelSHAP incur substantial higher computational overhead, due to internal optimization processes. Overall, our method achieves both faster convergence and improved computational efficiency, which demonstrates the benefits of directly estimating the group value rather than aggregating individual Shapley estimates.
@@ 4.2 Application to faithful copyright attribution in generative AI
Group data valuation is important for fairly compensating copyright holders whose data are used to train generative AI models. Existing approaches, such as the Shapley Royalty Share (SRS) proposed by Wang et al. (2024), adopt the GaI approach based on GSV. Given a partition of the training data into $K$ disjoint groups $S_{1},\dots,S_{K}$, SRS is defined as $\operatorname{SRS}(S_{k};x_{\mathrm{gen}}):=\frac{\operatorname{GSV}(S_{k};x_{\mathrm{gen}})}{\sum_{j=1}^{K}\operatorname{GSV}(S_{j};x_{\mathrm{gen}})}$. However, as discussed in Section 2, GSV is prone to the shell company attack. To address this vulnerability, we propose the Faithful Shapley Royalty Share (FSRS) that replaces GSV with our $\operatorname{FGSV}$ to faithfully reward individuals in a group by their contributions, not by tactical grouping strategies:
@@
Following Wang et al. (2024), we fine-tune Stable Diffusion v1.4 (Rombach et al., 2022) using Low-Rank Adaptation (LoRA; Hu et al. (2022)) on four brand logos from FlickrLogo-27 (Kalantidis et al., 2011). The utility $U(\cdot;x^{(\mathrm{gen})})$ is the average log-likelihood of generating $20$ brand-specific images $x^{(\mathrm{gen})}$ using the prompt “A logo by [brand name]” (see example images in Panel (a) of Figure 3). We compare SRS and FSRS under two grouping scenarios: (1) $30$ images from each brand form a single group, and (2) the Google and Sprite datasets are each split into two subgroups (20/10 images), launching a shell company attack. Importantly, the total data per brand remains unchanged across two scenarios.
Figure 3: Comparison of SRS and FSRS for copyright attribution. (a) Example images generated using brand prompts. (b) Shapley Royalty Share (SRS, (Wang et al., 2024)) based on GSV. (c) Faithful SRS (FSRS, our method) based on $\operatorname{FGSV}$. Blue bars: valuation under Scenario 1 (single group per brand); orange bars: valuation under Scenario 2 (Google/Sprite data each split into size-20/10 subgroups, colored in dark and light orange).
@@ 4.3 Application to faithful explainable AI
Group data valuation is also a crucial tool in explainable AI, providing interpretable summaries of data contributions at the group level— particularly in contexts where groups correspond to socially or scientifically significant categories. For example, Ghorbani and Zou (2019) used GSV to quantify the contributions of data from different demographic groups to patient readmission prediction accuracy (Strack et al., 2014). However, GSV’s sensitivity to the choice of grouping can cause group values to fluctuate dramatically under different data partitions, leading to inconsistent interpretations. To address this problem, we recommend practitioners to use FGSV and empirically demonstrate that it produces more consistent and reliable interpretations across a variety of grouping configurations.
@@
We conduct our experiment on the Diabetes dataset (Efron et al., 2004), which contains 442 individuals, each described by 10 demographic and health-related features (e.g., sex, age, and BMI). The task is to predict the progression of diabetes one year after baseline. We construct 7 grouping schemes, partitioned by all non-empty subsets of the variables sex, age, and BMI, excluding the trivial no-partition case. For the continuous variables age and BMI, we discretize each into three quantile-based categories. Overall, we will have between 2 to 18 total groups. For each grouping scheme, we compute GSV exactly and estimate FGSV via 30 Monte Carlo replications. Our predictive model is ridge regression, and we measure utility as the negative mean squared error on a held-out test set, with the null utility set to the variance of the test responses. Thus, higher group valuations indicate greater contributions to the model’s predictive accuracy. To compare category-level contributions across different grouping schemes, we aggregate group values as follows: for a given variable (e.g., sex), we sum the (F)GSVs of all groups that include that category. This allows for consistent cross-scheme comparisons, where one would expect a category to receive stable valuations after aggregation across different grouping strategies if the valuation method is robust.
Figure 4: Comparison of GSV (top row) and FGSV (bottom row) in a regression task for explainable AI. Each column aggregates category-level values for a specific variable: sex (left), age (middle), and BMI (right). Shaded areas represent $\pm 1$ standard deviation across 30 replications.
@@ 5 Conclusion and discussion
Beyond the shell company attack, there exists another unfair competition strategy, namely, the copier attack, in which a group may steal valuation from other groups by duplicating their high-value data points. An ad-hoc remedy is to pre-process the dataset to detect and remove such duplicated entries before applying our FGSV method. In existing literature, Falconer et al. (2025) suggests designing the utility function using Pearl’s “do operation”. This defense can be incorporated into our framework, letting our method also defend against the copier attack; but such do-utility function is not always available. Han et al. (2022) shows that with a submodular utility function, Banzhaf (Wang and Jia, 2023b) and leave-one-out can successfully discourage a genuine contributor from duplicating itself to gain higher total valuation; however, their method is prone to “pure infringers” who only copy from valuable data points without contributing original contents. Overall, it remains an open challenge to defend against the copier attack for general utility functions, while maintaining the defense against the shell company attack to ensure a safe group data valuation.